# Learning to Play Two-Player Perfect-Information Games without Knowledge

### v4

@@ 3.3 Comparison of Learning Data Selection Algorithms
We now compare tree learning, root learning and terminal learning, using the protocol of Section 3.2. Each combination uses either tree learning or root learning or terminal learning. Moreover, each combination uses either iterative deepening alpha-beta (denoted by $\mathrm{ID}$) or MCTS. Furthermore, each combination uses $\epsilon$-greedy as action selection method (see Section 3.1) and the classical terminal evaluation ($1$ if the first player wins, $-1$ if the first player loses, 00 in case of a draw). There are a total of $6$ combinations. The experiment was repeated $32$ times. The winning percentage of a combination for each game and for each evaluation step (i.e. each hour) is therefore calculated from $192$ matches. The winning percentage curves are shown in Figure 2. The final winning percentages are shown in Table 3. Each percentage of the table has required $6144$ matches. $\mathrm{ID}$ is first on all games except in Outer Open Gomoku where it is second (MCTS root learning is first) and in Surakarta (MCTS with tree learning is first). MCTS with root learning is better than MCTS with tree learning except in Breakthrough and Surakarta. At Hex and Amazons, MCTS with root learning gives better results throughout the learning process but ends up being caught up by $\mathrm{ID}$ with terminal learning. Terminal learning performs worse everywhere, except in a few cases where it is very slightly better. On average, $\mathrm{ID}$ with tree learning is better ($71\%$ win), then MCTS with root learning is second ($9\%$ lower win percentage), followed by MCTS with tree learning ($18\%$ lower to $\mathrm{ID}$).
@@
Figure 2: Evolutions of the winning percentages of the combinations of the experiment of Section 3.3, i.e. MCTS (dotted line) or iterative deepening alpha-beta (continuous line) with tree learning (blue line) or root learning (red line) or terminal learning (green line). The display uses a simple moving average of 6 data.
@@
Table 3: Final winning percentages of the combinations of the experiment of Section 3.3 ($\mathrm{ID}$: iterative deepening alpha-beta). Reminder: the percentage is the average over the repetitions, of the winning percentage of a combination against each other combination, in first and second player (see 3.2 ; $95\%$ confidence intervals: max ±$\pm 0.85\%$).
@@ 4 Tree Search Algorithms for Game Learning
In this section, we introduce a new tree search algorithm, that we call descent minimax or more succinctly descent, dedicated to be used during the learning process. It requires tree learning (combining it with root learning or terminal learning is of no interest [^6] Because, with root and terminal learning, the extra computations of descent compared to $\mathrm{UBFM}$ with the same number of iterations, do not change, in general, the values to learn and the move to play (but the time to perform the iterations is much longer).). After presenting descent, we compare it to MCTS with root learning and with tree learning, to iterative deepening alpha-beta with root learning and with tree learning and to $\mathrm{UBFM}$ with tree learning.
@@ 4.1 Descent: Generate Better Data
                              $S\leftarrow S\cup\{a(s)\}$
                              $v^{\prime}(s,a)\leftarrow f_{\mathrm{t}}\left(a(s)\right)$
                              $v(a(s))\leftarrow v^{\prime}(s,a)$
@@
                              $v^{\prime}(s,a)\leftarrow f_{\theta}\left(a(s)\right)$
            $a_{b}\leftarrow$ best_action($s$)
            $v^{\prime}(s,a_{b})\leftarrow$ descent_iteration($a_{b}(s)$)
            $a_{b}\leftarrow$ best_action($s$)
            $v(s)\leftarrow v^{\prime}(s,a_{b})$
@@
            return ${\displaystyle\operatorname*{\mathrm{arg\,max}}_{a\in\textnormal{{actions(}}\textnormal{\emph{s}}\textnormal{{)}}}v^{\prime}\left(s,a\right)}$
@@
            return ${\displaystyle\operatorname*{\mathrm{arg\,min}}_{a\in\textnormal{{actions(}}\textnormal{\emph{s}}\textnormal{{)}}}v^{\prime}\left(s,a\right)}$
@@
Algorithm 8 Descent tree search algorithm (see Table 1 for the definitions of symbols ; note: $S$ is the set of states which are non-leaves or terminal and $T=(v,v^{\prime})$).
Thus, we present descent. It is a modification of $\mathrm{UBFM}$ which builds a different, deeper, game tree, to be combined with tree learning. The idea of descent is to combine $\mathrm{UBFM}$ with deterministic end-game simulations providing interesting values from the point of view of learning. The algorithm descent (Algorithm 8) recursively selects the best child of the current node, which becomes the new current node. It adds the children of the current node if they are not in the tree. It performs this recursion from the root (the current state of the game) until reaching a terminal node (an end game). It then updates the value of the selected nodes (minimax value). The algorithm descent repeats this recursive operation starting from the root as long as there is some search time left. Descent is almost identical to $\mathrm{UBFM}$. The only difference is that descent performs an iteration until reaching a terminal state while $\mathrm{UBFM}$ performs this iteration until reaching a leaf of the tree ($\mathrm{UBFM}$ stops the iteration much earlier). In other words, during an iteration, $\mathrm{UBFM}$ just extends one of the leaves of the game tree while descent recursively extends the best child from this leaf until reaching the end of the game. The algorithm descent has the advantage of $\mathrm{UBFM}$, i.e. to perform a longer search to determine a better action to play. By learning the values of the game tree (by using for example tree learning), it also has the advantage of a minimax search at depth $1$, i.e. to raise the values of the terminal nodes to the other nodes more quickly. In addition, the states thus generated are closer to the terminal states. Their values are therefore better approximations.
@@ 4.2 Comparison of Search Algorithms for Game Learning
We now compare descent with tree learning to MCTS with root learning and with tree learning, to iterative deepening alpha-beta with root learning and with tree learning, and to $\mathrm{UBFM}$ with tree learning, using the protocol of Section 3.2. Each combination uses one of these tree search algorithms combined with tree/root learning. There are a total of $6$ combinations. The experiment was repeated $32$ times. The winning percentage of a combination for each game and for each evaluation step (i.e. each hour) is therefore calculated from $192$ matches. The winning percentage curves are shown in Figure 3. The final winning percentages are shown in Table 4. Each percentage of the table has required $6144$ matches. It is descent which gets the best curves on all games. For two games (Surakarta and Outer Open Gomoku), the difference with $\mathrm{UBFM}$ is very narrow but the results remain better than the classic approaches (MCTS and alpha-beta). On each game, descent obtains a final percentage higher than all the other combinations (except in Santorini where it is $2\%$ lower than $\mathrm{UBFM}$, the best algorithm at this game). On average over all games, descent has $82\%$ win and is above $\mathrm{UBFM}$, the second best combination, by $18\%$ and $\mathrm{ID}$ with tree learning, the third best combination, by $34\%$.
Figure 3: Evolutions of the winning percentages of the combinations of the experiment of Section 4.2, i.e. of descent (dashed line), $\mathrm{UBFM}$ (dotted dashed line), MCTS (dotted line), iterative deepening alpha-beta (continuous line) with tree learning (blue line) or root learning (red line). The display uses a simple moving average of 6 data.
@@
Table 4: Final winning percentages of the combinations of the experiment of Section 4.2 ($\mathrm{ID}$: iterative deepening alpha-beta ; see 3.2 ; $95\%$ confidence intervals: max ±$\pm 0.85\%$)
In conclusion, descent (with tree learning) is undoubtedly the best combination. $\mathrm{UBFM}$ (with tree learning) is the second best combination, sometimes very close to descent performances and sometimes very far, but always superior to other combinations (slightly or largely depending on the game).
5 Reinforcement Heuristic to Improve Learning Performance
In this section, we propose the technique of reinforcement heuristic, which consists to replace the classical terminal evaluation function – that we denote by $\mathrm{b_{t}}$, which returns $1$ if the first player wins, $-1$ if the second player wins, and 00 in case of a draw [36, 2, 16] – by another heuristic to evaluate terminal states during the learning process. By using this technique, non-terminal states are therefore evaluated differently, partial game trees and thus matches during the learning process are different, which can impact the learning performances. We start by defining what we call a reinforcement heuristic and we offering several reinforcement heuristics. Then, we propose a complementary technique, that we call completion, which corrects state evaluation functions taking into account the resolution of states. Finally, we compare the reinforcement heuristics that we propose to the classical terminal evaluation function.
5.1 Some Reinforcement Heuristics
Definition 6.
5.1.1 Scoring
Some games have a natural reinforcement heuristic: the game score. For example, in the case of the game Othello (and in the case of the game Surakarta), the game score is the number of its pieces minus the number of pieces of his opponent (the goal of the game is to have more pieces than its opponent at the end of the game). The scoring heuristic used as a reinforcement heuristic consists of evaluating the terminal states by the final score of the game. With this reinforcement heuristic, the adaptive evaluation function will seek to learn the score of states. In the context of an algorithm based on minimax, the score of a non-terminal state is the minimax value of the subtree starting from this state whose terminal leaves are evaluated by their scores. After training, the adaptive evaluation function then contains more information than just an approximation of the result of the game, it contains an approximation of the score of the game. If the game score is intuitive, this should improve learning performances.
@@ 5.1.1 Scoring
In the context of the game of the Amazons, the score is the size of the territory of the winning player, i.e. the squares which can be reached by a piece of the winning player. This is approximately the number of empty squares.
5.1.2 Additive and Multiplicative Depth Heuristics
Now we offer the following reinforcement heuristic: the depth heuristic. It consists in giving a better value to the winning states close to the start of the game than to the winning states far from the start. Reinforcement learning with the depth heuristic, it is learning the duration of matches in addition to their results. This learned information is then used to try to win as quickly as possible and try to lose as late as possible. The hypothesis of this heuristic is that a state close to the end of the game has a more precise value than a state more distant and that the duration of the game is easily learned. Under this assumption, with this heuristic, we will take less risk to try to win as quickly as possible and to lose as late as possible. In addition, with a long game, a player in difficulty will have more opportunities to regain the upper hand. We propose two realizations of the depth heuristic: the additive depth heuristic, that we denote by $\mathrm{p_{t}}$, and the multiplicative depth heuristic, that we denote by $\mathrm{p_{t}}^{\prime}$. The evaluation function $\mathrm{p_{t}}$ returns the value $l$ if the first player wins, the value $-l$ if the second player wins, and 00 in case of a draw, with $l=P-p+1$ where $P$ is the maximum number of playable actions in a game and $p$ is the number of actions played since the beginning of the game. For the game of Hex, $l$ is the number of empty cells on the board plus $1$. For the games where $P$ is very large or difficult to compute, we can instead use $l=\max\left(1,\tilde{P}-p\right)$ with $\tilde{P}$ a constant approximating $P$ (close to the empirical average length of matches).The evaluation function $\mathrm{p_{t}}^{\prime}$ is identical except that $l$ satisfies $l=\frac{\tilde{P}}{p}$.
Remark 8.
5.1.3 Cummulative Mobility
The next reinforcement heuristic that we propose is cummulative mobility. It consists in favoring the games where the player has more possibility of action and where his opponent has less. The implementation used in this article is as following. The value of a terminal state is $\frac{M_{1}}{M_{2}}$ if the first player wins, $-\frac{M_{2}}{M_{1}}$ if the second player wins, and 00 in case of a draw, where $M_{1}$ is the mean of the number of available actions in each turn of the first player since the start of the game and $M_{2}$ is the mean of the number of available actions in each turn of the second player since the start of the game.
5.1.4 Piece Counting: Presence
Finally, we propose as reinforcement heuristic: the presence heuristic. It consists in taking into account the number of pieces of each player and starts from the assumption that the more a player has pieces the more this one has an advantage. There are several implementations for this heuristic, we use in this article the following implementation: the heuristic value is $\max(n_{1}-n_{2},1)$ if the first player wins, $\min(n_{1}-n_{2},-1)$ if the second player wins, and 00 in case of a draw, where $n_{1}$ is the number of pieces of the first player and $n_{2}$ is the number of pieces of the second player. Note that in the games Surakarta and Othello, the score corresponds to a presence heuristic.
5.2 Completion
Relying solely on the value of states calculated from the terminal evaluation function and the adaptive evaluation function can sometimes lead to certain aberrant behaviors. More precisely, if we only seek to maximize the value of states, we will then choose to play a state $s$ rather than another state $s^{\prime}$ if $s$ is of greater value than $s^{\prime}$ even if $s^{\prime}$ is a winning resolved state (a state is resolved if we know the result of the game starting from this state in which the two players play optimally). A search algorithm can resolve a state. This happens when all the leaves of the subtree starting from this state are terminal. Choosing $s$ rather than $s^{\prime}$, a winning resolved state, is an error [^7] There is perhaps, in certain circumstances, an interest in making this error from the point of view of learning. when $s$ is not resolved (or when $s$ is resolved and is not winning). By choosing $s$, guarantee of winning is lost. The left graph of Figure 4 illustrates such a scenario.
@@ 5.2 Completion
It is therefore necessary to take into account both the value of states and the resolution of states. The completion technique, which we propose in this section, is one way of doing it. It consists, on the one hand, in associating with each state $s$ a completion value $c(s)$ and a resolution value $r(s)$. The completion value $c(s)$ of a leaf state $s$ is 00 if the state $s$ is not terminal or if it is a draw, $1$ if it is a winning terminal state and $-1$ if the state is a losing terminal state. The value $c(s)$ of a non-leaf state $s$ is computed as the minimax value of the subtree of the partial game tree starting from $s$ where the leaves are evaluated by their completion value. The resolution value $r(s)$ of a leaf state $s$ is 00 if the state $s$ is not terminal and $1$ if it is terminal. The resolution value $r(s)$ of a non-leaf state $s$ is $1$ if $\left|c\left(s\right)\right|=1$. Otherwise, $r(s)$ is the minimum of the resolution values of the childen of $s$. The completion technique consists, on the other hand, in using $c\left(\cdot\right)$ to compute $v\left(\cdot\right)$. For this, states are compared from pairs $\left(c\left(\cdot\right),v\left(\cdot\right)\right)$, by using the lexicographic order (instead of just compare states from $v\left(\cdot\right)$). More precisely, the value $v\left(s\right)$ of a state $s$ is computed in calculating $\left(c(s),v(s)\right)$ as the minimax value of the subtree of the partial game tree starting from $s$ where the leaves $l$ are evaluated by $\left(c(l),v(l)\right)$. Thus, the value $v\left(s\right)$ of a winning state $s$ is always the value of the corresponding winning terminal leaf. Finally, with the completion technique, to decide which action to play during the search, we choose the best unresolved action if it exists and otherwise the best resolved action (i.e. we choose the action which maximize $\left(-r\left(s\right),c\left(s\right),v\left(s\right)\right)$ in a max state and which minimize $\left(r\left(s\right),c\left(s\right),v\left(s\right)\right)$ in a min state). The right graph of Figure 4 illustrates the use of completion. The use of the resolution of states also makes it possible to stop the search in the resolved subtrees and thus to save computing time. Descent algorithm modified to use the completion and the resolution stop is described in Algorithm 9. With completion, descent always chooses an action leading to a winning resolved state and never chooses, if possible, an action leading to a losing resolved state.
Remark 9.
Remark 10.
A variant consists in performing a strong resolution, i.e. to compute $r\left(\cdot\right)$ for non-leaf states in the following way: $r\left(s\right)$ is (in all cases) the minimum of the resolution values of the childen of the state $s$. With this variant, unnecessary calculations are made to determine the best action to play. However, the additional values may be useful for learning. It is not clear which variant is the best.
Another variant is possible, it consists in not using the completition $c\left(\cdot\right)$ to compute $v\left(\cdot\right)$. With this version, $v\left(s\right)$ is the minimax value of the subtree starting from $s$ where the leaf nodes $l$ are evaluated by $v\left(l\right)$. With this variant, for certain resolved states, instead of learning a lower bound (in absolute value), which is an exact value, we learn an estimated value, which could be a potentially an overestimated value. It is not clear which variant is the best.
Function completed_descent_iteration($s$, $S$, $T$, $f_{\theta}$, $f_{\mathrm{t}}$)
@@
                        if terminal($a(s)$) then
                              $S\leftarrow S\cup\{a(s)\}$
                              $v^{\prime}(s,a)\leftarrow f_{\mathrm{t}}\left(a(s)\right)$
                              $r(a(s)),c(a(s)),v(a(s))\leftarrow 1,\mathrm{b_{t}}(a(s)),v^{\prime}(s,a)$
@@
                              $v^{\prime}(s,a)\leftarrow f_{\theta}\left(a(s)\right)$
@@
                  $c(s),v(s)\leftarrow c\left(a_{b}(s)\right),v^{\prime}\left(s,a_{b}\right)$
@@
                  $A\leftarrow\left\{a\in\textnormal{{actions(}}\textnormal{\emph{s}}\textnormal{{)}}\,|\,r\left(a\left(s\right)\right)=0\right\}$
                  $a_{b}\leftarrow$ completed_best_action_dual($s$, $A$)
                  $n(s,a_{b})\leftarrow n(s,a_{b})+1$
                  $v^{\prime}\left(s,a_{b}\right)\leftarrow$ completed_descent_iteration($a_{b}(s)$)
                  $a_{b}\leftarrow$ completed_best_action($s$, actions(s))
                  $c(s),v(s)\leftarrow c\left(a_{b}(s)\right),v^{\prime}\left(s,a_{b}\right)$
@@
      return $v\left(s\right)$
@@
      while time()$-\,t<\tau$wedge $r\left(s\right)=0$ do completed_descent_iteration($s$, $S$, $T$, $f_{\theta}$, $f_{\mathrm{t}}$)
@@
Algorithm 9 Descent tree search algorithm with completion and resolution stop (see Table 1 for the definitions of symbols and Algorithm 10 for the definitions of completed_best_action($s$) and backup_resolution($s$)). Note: $T=(v,v^{\prime},c,r)$ and $S$ is the set of states of the partial game tree which are non-leaves or terminal.
@@
            return ${\displaystyle\operatorname*{\mathrm{arg\,max}}_{a\in A}\left(c\left(a\left(s\right)\right),v^{\prime}\left(s,a\right),n\left(s,s^{\prime}\right)\right)}$
@@
            return ${\displaystyle\operatorname*{\mathrm{arg\,min}}_{a\in A}\left(c\left(a\left(s\right)\right),v^{\prime}\left(s,a\right),-n\left(s,s^{\prime}\right)\right)}$
@@
            return ${\displaystyle\operatorname*{\mathrm{arg\,max}}_{a\in A}\left(c\left(a\left(s\right)\right),v^{\prime}\left(s,a\right),-n\left(s,s^{\prime}\right)\right)}$
@@
            return ${\displaystyle\operatorname*{\mathrm{arg\,min}}_{a\in A}\left(c\left(a\left(s\right)\right),v^{\prime}\left(s,a\right),n\left(s,s^{\prime}\right)\right)}$
@@
            return $\operatorname*{\mathrm{min}}{}_{a\in\textnormal{{actions(}}\textnormal{\emph{s}}\textnormal{{)}}}r\left(a\left(s\right)\right)$
@@
We also propose to use the resolution of states with action selections, to reduce the duration of games and therefore a priori the duration of the learning process: always play an action leading to a winning resolved state if it exists and never play an action leading to a losing resolved state if possible. Thus, if among the available actions we know that one of the actions is winning, we play it. If there is none, we play according to the chosen action selection method among the actions not leading to a resolved state (if possible). If it is not possible, we play the best action leading to a state resolved as a draw and if there is none, we play the best action leading to a losing resolved state. We call it completed action selection.

### v5

@@ 3.4 Comparison of Learning Data Selection Algorithms
We now compare tree learning, root learning and terminal learning, using the protocol of Section 3.3. Each combination uses either tree learning, root learning, or terminal learning. Moreover, each combination uses either iterative deepening alpha-beta (denoted by $\mathrm{ID}$) or MCTS. Furthermore, each combination uses $\epsilon$-greedy as action selection method (see Section 3.1) and the classical terminal evaluation ($1$ if the first player wins, $-1$ if the first player loses, and 00 in case of a draw). There are a total of $6$ combinations. The experiment was repeated $32$ times. The winning percentage of a combination for each game and for each evaluation step (i.e. each hour) is therefore calculated from $384$ matches (recall that there is a first player match and a second player match per pair of combinations).
The winning percentage curves are shown in Figure 1. The final winning percentages are shown in Table 2. Each percentage of the table has required $12,288$ matches. $\mathrm{ID}$ is first on all games except in Outer Open Gomoku where it is second (MCTS root learning is first) and in Surakarta (MCTS with tree learning is first). MCTS with root learning is better than MCTS with tree learning except in Breakthrough and Surakarta. At Hex and Amazons, MCTS with root learning gives better results throughout the learning process but ends up being caught up by $\mathrm{ID}$ with terminal learning. Terminal learning performs worse everywhere, except in a few cases where it is very slightly better. On average, $\mathrm{ID}$ with tree learning is better ($71\%$ win), then MCTS with root learning is second ($9\%$ lower win percentage), followed by MCTS with tree learning ($18\%$ lower to $\mathrm{ID}$).
@@
Figure 1: Evolutions of the winning percentages of the combinations of the experiment of Section 3.4, i.e. MCTS (dotted line) or iterative deepening alpha-beta (continuous line) with tree learning (blue line), root learning (red line), or terminal learning (green line). The display uses a simple moving average of 6 data.
@@
Table 2: Final winning percentages of the combinations of the experiment of Section 3.4 ($\mathrm{ID}$: iterative deepening alpha-beta). Reminder: the percentage is the average over the repetitions, of the winning percentage of a combination against each other combination, in first and second player (see 3.3 ; $95\%$ confidence intervals: max ±$\pm 0.85\%$).
@@ 4 Tree Search Algorithms for Game Learning
In this section, we introduce a new tree search algorithm, that we call Descent Minimax or more succinctly Descent, dedicated to be used during the learning process. After presenting Descent, we compare it to MCTS with root learning and with tree learning, to iterative deepening alpha-beta with root learning and with tree learning and to $\mathrm{UBFM}$ with tree learning.
@@ 4.1 Descent: Generate Better Data
                              $v(s,a)\leftarrow$ descent_iteration($a(s),S,T,f_{\theta},f_{\mathrm{t}}$)
@@
                              $v(s,a)\leftarrow f_{\theta}\left(a(s)\right)$
            $a\leftarrow$ best_action($s$)
            $v(s,a)\leftarrow$ descent_iteration($a(s),S,T,f_{\theta},f_{\mathrm{t}}$)
            $a\leftarrow$ best_action($s$)
            $v(s)\leftarrow v(s,a)$
@@
            return ${\displaystyle\operatorname*{\mathrm{arg\,max}}_{a\in\textnormal{{actions(}}\textnormal{\emph{s}}\textnormal{{)}}}v\left(s,a\right)}$
@@
            return ${\displaystyle\operatorname*{\mathrm{arg\,min}}_{a\in\textnormal{{actions(}}\textnormal{\emph{s}}\textnormal{{)}}}v\left(s,a\right)}$
@@
Algorithm 9 Descent minimax algorithm (see Table 1 for the definitions of symbols ; note: $S$ is the set of states which are non-leaves or terminal and $T=(v,)$).
Thus, we present Descent. It is a modification of $\mathrm{UBFM}$ which builds a different, deeper, game tree, to be combined with tree learning.
Remark 6.
Note that, combining it with root learning or terminal learning is of no interest. The reason is that Descent spend most of its time exploring a part of the game tree whose only purpose is to provide states for learning. With root or terminal learning, these states are not learned and therefore generated for nothing.
The idea of Descent is to combine $\mathrm{UBFM}$ with deterministic end-game simulations providing interesting values from the point of view of learning.
The Descent algorithm (Algorithm 9) recursively selects the best child of the current node, which becomes the new current node. It performs this recursion from the root (the current state of the game) until reaching a terminal node (an end game). It then updates the value of the selected nodes (minimax value). The Descent algorithm repeats this recursive operation starting from the root as long as there is some search time left. Descent is almost identical to $\mathrm{UBFM}$. The only difference is that Descent performs an iteration until reaching a terminal state while $\mathrm{UBFM}$ performs this iteration until reaching a leaf of the tree ($\mathrm{UBFM}$ stops the iteration much earlier). In other words, during an iteration, $\mathrm{UBFM}$ just extends one of the leaves of the game tree while Descent recursively extends the best child from this leaf until reaching the end of the game.
The Descent algorithm has the advantage of $\mathrm{UBFM}$, i.e. to perform a longer search to determine a better action to play. By learning the values of the game tree (by using for example tree learning), it also has the advantage of a minimax search at depth $1$, i.e. to raise the values of the terminal nodes to the other nodes more quickly. In addition, the states thus generated are closer to the terminal states. Their values are therefore better approximations.
@@ 4.2 Comparison of Search Algorithms for Game Learning
We now compare Descent with tree learning to MCTS with root learning and with tree learning, to iterative deepening alpha-beta with root learning and with tree learning, and to $\mathrm{UBFM}$ with tree learning, using the protocol of Section 3.3. There are a total of $6$ combinations. The experiment was repeated $32$ times. The winning percentage of a combination for each game and for each evaluation step (i.e. each hour) is therefore calculated from $384$ matches (recall that there is a first player match and a second player match per pair of combinations).
The winning percentage curves are shown in Figure 2. The final winning percentages are shown in Table 3. Each percentage of the table has required $12,288$ matches. It is Descent which gets the best curves on all games. For two games (Surakarta and Outer Open Gomoku), the difference with $\mathrm{UBFM}$ is very narrow but the results remain better than the classic approaches (MCTS and alpha-beta). On each game, Descent obtains a final percentage higher than all the other combinations (except in Santorini where it is $2\%$ lower than $\mathrm{UBFM}$, the best algorithm at this game). On average over all games, Descent has $82\%$ win and is above $\mathrm{UBFM}$, the second best combination, by $18\%$. It is also above $\mathrm{ID}$ with tree learning, the third best combination, by $34\%$.
Figure 2: Evolutions of the winning percentages of the combinations of the experiment of Section 4.2, i.e. of Descent (dashed line), $\mathrm{UBFM}$ (dotted dashed line), MCTS (dotted line), and iterative deepening alpha-beta (continuous line) with tree learning (blue line) or root learning (red line). The display uses a simple moving average of 6 data.
@@
Table 3: Final winning percentages of the combinations of the experiment of Section 4.2 ($\mathrm{ID}$: iterative deepening alpha-beta ; see 3.3 ; $95\%$ confidence intervals: max ±$\pm 0.85\%$)
In conclusion, Descent is the best search algorithm for learning evaluation functions. $\mathrm{UBFM}$ (with tree learning) is the second best algorithm, sometimes very close to Descent performances and sometimes very far, but always superior to the other algorithms (slightly or largely depending on the game).
5 Completion
In this section, we propose a complementary technique, that we call completion, which corrects state evaluation functions taking into account the resolution of states.
@@ 5 Completion
However, we do not evaluate this technique in this paper, to avoid making this paper unnecessarily cumbersome and to focus on the learning aspects that constitute the main contributions of this paper. The evaluation of this technique is however available in [35] which studies the impact of improvements of Unbounded Minimax. This study shows that completion improves the performance of Unbounded Minimax by $6\%\pm 0.7\%$.
5.1 Completion Concept
Relying solely on the value of states calculated from the terminal evaluation function and the adaptive evaluation function can sometimes lead to certain aberrant behaviors. More precisely, if we only seek to maximize the value of states, we will then choose to play a state $s$ rather than another state $s^{\prime}$ when $s$ is of greater value than $s^{\prime}$ even if $s^{\prime}$ is a winning resolved state (a state is resolved if we know the result of the match starting from that state in which the two players play optimally). A search algorithm can resolve a state. This happens when all the leaves of the subtree starting from that state are terminal. Choosing $s$ rather than $s^{\prime}$, a winning resolved state, is an error [^6] There is perhaps, in certain circumstances, an interest in making this error from the point of view of learning. when $s$ is not resolved (or when $s$ is resolved and is not winning). By choosing $s$, guarantee of winning is lost. The left graph of Figure 3 illustrates such a scenario.
@@ 5.1 Completion Concept
It is therefore necessary to take into account both the value of states and the resolution of states.
5.2 Completed Algorithms
@@
The completion technique, which we propose in this section, is one way of doing it. It consists, on the one hand, in associating with each state $s$ a completion value $c(s)$ and a resolution value $r(s)$. The completion value $c(s)$ of a leaf state $s$ is 00 if the state $s$ is not terminal or if it is a draw, $1$ if it is a winning terminal state, and $-1$ if the state is a losing terminal state. The value $c(s)$ of a non-leaf state $s$ is computed as the minimax value of the subtree of the partial game tree starting from $s$ where the leaves are evaluated by their completion value. The resolution value $r(s)$ of a leaf state $s$ is 00 if the state $s$ is not terminal and $1$ if it is terminal. The resolution value $r(s)$ of a non-leaf state $s$ is $1$ if $\left|c\left(s\right)\right|=1$. Otherwise, $r(s)$ is the minimum of the resolution values of the childen of $s$.
The completion technique consists, on the other hand, in using $c\left(\cdot\right)$ to compute $v\left(\cdot\right)$. For this, states are compared from pairs $\left(c\left(\cdot\right),v\left(\cdot\right)\right)$, by using the lexicographic order (instead of just compare states from $v\left(\cdot\right)$). More precisely, the value $v\left(s\right)$ of a state $s$ is computed in calculating $\left(c(s),v(s)\right)$ as the minimax value of the subtree of the partial game tree starting from $s$ where the leaves $l$ are evaluated by $\left(c(l),v(l)\right)$. Thus, the value $v\left(s\right)$ of a winning state $s$ is always the value of the corresponding winning terminal leaf. The right graph of Figure 3 illustrates the use of completion.
Finally, with the completion technique, to decide which action to play during the search, we choose the best unresolved action if it exists and otherwise the best resolved action (i.e. we choose the action which maximize $\left(-r\left(s\right),c\left(s\right),v\left(s\right)\right)$ in a max state and which minimize $\left(r\left(s\right),c\left(s\right),v\left(s\right)\right)$ in a min state).
Without completion, algorithms such as Descent or Unbounded Minimax can get stuck in a non-optimal fixed point. The completion technique ensures that with sufficient thinking time, the optimal strategy is found [51].
Moreover, the use of the resolution of states also makes it possible to stop the search in the resolved subtrees and thus to save computing time. Descent algorithm modified to use the completion and the resolution stop is described in Algorithm 10. With completion, after the search, Unbounded Minimax always chooses an action leading to a winning resolved state and never chooses, if possible, an action leading to a losing resolved state. Unbounded Minimax algorithm modified to use the completion and the resolution stop is described in Algorithm 13.
Function descent_iteration($s$, $S$, $T$, $f_{\theta}$, $f_{\mathrm{t}}$)
@@
                        if terminal($a\left(s\right)$) then
                              $r(s,a),c(s,a),v(s,a)\leftarrow$ descent_iteration($a(s),S,T,f_{\theta},f_{\mathrm{t}}$)
@@
                              $v(s,a)\leftarrow f_{\theta}\left(a(s)\right)$
@@
                  $c(s),v(s)\leftarrow c\left(s,a_{b}\right),v\left(s,a_{b}\right)$
@@
                  $A\leftarrow\left\{a\in\textnormal{{actions(}}\textnormal{\emph{s}}\textnormal{{)}}\,|\,r\left(s,a\right)=0\right\}$
                  $a\leftarrow$ completed_best_action_dual($s$, $A$)
                  $n(s,a)\leftarrow n(s,a)+1$
                  $r\left(s,a\right),c\left(s,a\right),v\left(s,a\right)\leftarrow$ descent_iteration($a(s),S,T,f_{\theta},f_{\mathrm{t}}$)
                  $a\leftarrow$ completed_best_action($s$, actions(s))
                  $c(s),v(s)\leftarrow c\left(s,a\right),v\left(s,a\right)$
@@
      return $r\left(s\right),c\left(s\right),v\left(s\right)$
@@
      while time()$-\,t<\tau$wedge $r\left(s\right)=0$ do descent_iteration($s$, $S$, $T$, $f_{\theta}$, $f_{\mathrm{t}}$)
@@
Algorithm 10 Descent tree search algorithm with completion and resolution stop (see Table 1 for the definitions of symbols and Algorithm 11 for the definitions of completed_best_action($s$) and backup_resolution($s$)). Note: $T=(v,c,r)$ and $S$ is the set of states of the partial game tree which are non-leaves or terminal.
@@
            return ${\displaystyle\operatorname*{\mathrm{arg\,max}}_{a\in A}\left(c\left(s,a\right),v\left(s,a\right),n\left(s,s^{\prime}\right)\right)}$
@@
            return ${\displaystyle\operatorname*{\mathrm{arg\,min}}_{a\in A}\left(c\left(s,a\right),v\left(s,a\right),-n\left(s,s^{\prime}\right)\right)}$
@@
            return ${\displaystyle\operatorname*{\mathrm{arg\,max}}_{a\in A}\left(c\left(s,a\right),v\left(s,a\right),-n\left(s,s^{\prime}\right)\right)}$
@@
            return ${\displaystyle\operatorname*{\mathrm{arg\,min}}_{a\in A}\left(c\left(s,a\right),v\left(s,a\right),n\left(s,s^{\prime}\right)\right)}$
@@
            return $\operatorname*{\mathrm{min}}{}_{a\in\textnormal{{actions(}}\textnormal{\emph{s}}\textnormal{{)}}}r\left(s,a\right)$
@@
We also propose to use the resolution of states with action selections, to reduce the duration of games and therefore a priori the duration of the learning process: always play an action leading to a winning resolved state if it exists and never play an action leading to a losing resolved state if possible. Thus, if among the available actions we know that one of the actions is winning, we play it. If there is none, we play according to the chosen action selection method among the actions not leading to a losing resolved state (if possible). We call it completed action selection.